OpenStudy (epicteatime):

Steve can paint a room three times as fast as Billy. When they work together, Steve and Billy can paint a large room in 4 hours. How many hours would it take Billy to paint it by himself?

OpenStudy (epicteatime):

@ganeshie8

OpenStudy (epicteatime):

@Lyrae

ganeshie8 (ganeshie8):

\[\large \dfrac{1}{x} + \dfrac{3}{x} = \dfrac{1}{4}\] solve \(\large x\)